OpenStudy (anonymous):

Financial Math Question, will fan and medal! Please help really need it! Aaron needs $17,450 to purchase a motorcycle. How much money will hennaed to invest today in a savings account earning 3.8% interest, compounding monthly, to have enough money to purchase the boat in six yesars? A. $1,190.11 B. $11,075.97 C. $13,897.37 D. $17,056.70

OpenStudy (anonymous):

A=P(1+r/n)^nt A is the ending amount, P is the beginning amount, r is the interest rate (expressed as a decimal), n is the number of compoundings a year, t is the total number of years. If interest is compounded yearly, then n = 1; if semi-annually, then n = 2; quarterly, then n = 4; monthly, then n = 12; weekly, then n = 52; daily, then n = 365;

OpenStudy (anonymous):

solve for P P=A/(1+r/n)^nt

OpenStudy (anonymous):

A=17450 r=3.8/100 n=12 t=6